# Combining Neural Fields and Deformation Models for Non-Rigid 3D Motion Reconstruction from Partial Data

### v1

@@ 1 Introduction
Non-rigid 3D motion reconstruction involves recovering the shape and movement of objects undergoing arbitrary non-rigid motions based on visual observations. Given our naturally dynamic world, this task has extensive applications, particularly in digitizing natural scenes for virtual reality and entertainment. Our focus is on monocular depth observations, which can be easily captured using standard devices, including many consumer-level products.
@@
Figure 1: Given monocular depth observations of a moving shape, our approach produces complete reconstructions that preserve observed geometric details while establishing dense tracking. Our method is evaluated with motions of clothed humans (left) and animals (right).
@@
The first one encompasses parametric models, which use combined shape and motion parameters to handle specific entities (e.g. humans, faces, or animals). Examples include models like SCAPE [4], BlendSCAPE [23], and SMPL [35] for human figures, Flame [31] for faces, and SMAL [64] for animals. These parametric models have gained significant success, largely thanks to their ability to provide robust, temporally consistent estimations. However, they often lack generalizability across shape classes and struggle to capture geometric details outside of model constraints, such as hair or loose clothing in human representations.
@@
The second category includes methods that decouple shape and motion models, allowing for greater generalization across shape classes. Inspired by Niemeyer et al.[42], many approaches in this category use an implicit scene representation as a template, that undergoes unconstrained displacement fields (e.g., 3D flows) over time. While templates enforce consistency, they can restrict the model’s ability to depict finer geometric details. Additionally, the 3D flow-based models are often insufficiently constrained, leading to inconsistencies in motion representation. Another branch in this category focuses on time-based shape reconstructions rather than motion, either omitting explicit motion modeling altogether [59] or limiting it to frame pairs [32, 60]. Although useful for shape capture, these approaches lack broader applicability where motion is required and may yield topologically inconsistent reconstructions.
@@
Our approach combines an implicit neural field shape representation with a near-isometric mesh deformation model. This combination enables tracking a large class of 3D motions such as clothed 3D human motions and other vertebrate motions such as animals while benefiting from neural field representations for detailed reconstructions.
@@
To this end, we introduce a two-step data-driven approach. First, an encoder-decoder architecture equipped with an attention mechanism fuses input observations over a time sequence to infer full neural field reconstructions at each time step. Second, these reconstructions are fed into a deformation network that predicts inter-frame deformations by fitting the reconstructions to a near-isometric mesh deformation model. Both steps are trained simultaneously without motion supervision but with losses that promote geometric feature association between frames, near-isometric deformations, and 3D reconstruction losses. After training, per-frame reconstructions and their tracking are inferred from monocular observations in a single forward step.
@@
  • A representation of moving 3D shapes as neural signed distance fields, connected through a near-isometric surface deformation.
@@ 2 Related Works
Methods to reconstruct a possibly moving 3D shape can be categorized into two classes. On the one hand, array-based methods [7, 57, 54, 62, 63, 33, 20, 24] use multiple calibrated cameras that require costly setups and are thus restricted to professional use. On the other hand, depth-fusion-like methods [25, 21, 49, 58, 15] enable data obtained from commodity sensors to be used for consumer level applications. An active research direction aims to reconstruct possibly moving 3D shapes from sensor data, e.g. from a single RGB or RGB-D image [48, 22, 46], from an RGB video [3, 17, 1, 30, 10], from depth views [9, 59, 60, 56] or from point clouds [42, 50]. We review methods that input geometry observation, i.e. depth views or point clouds, as our method considers similar inputs. These approaches follow two main lines of works: model-based methods that leverage parametric shape models, and model-free methods that generalise to multiple shape classes.
@@ 2.1 Model-Based Methods
Early model-based strategies propose optimisation-based techniques. Weiss et al. [55] fit partial observations to the SCAPE [4] human body model. Mosh [34] uses a sparse set of markers to fit SCAPE [4] while Mosh++ [38] uses SMPL [35]. To augment the expressivity of parametric human models, several approaches [47, 2, 53] add vertex displacements on top of SMPL [35] to model clothes. More recently, data-driven approaches were proposed. IP-Net [6] combines parametric and implicit representations. H4D [26] proposes a compositional representation, which disentangles shape and motion. NSF [56] propose to combine SMPL [35] with a neural surface field to represent fine grained surface details. Neural Parametric Models (NPMs) [43] propose to learn custom disentangled shape and pose spaces from a dataset to which we can fit observations at inference. SPAMs [44] extend NPMs by learning disentangled semantic-part-based shape and pose spaces.
@@ 2.2 Model-Free Methods
Method Tracking Shape Completion Long Temp. Ctxt. Detail Preservation Unsupervised
@@
Some works consider static 3D reconstruction, e.g. Implicit Feature Networks (IF-Nets) [13]. IF-Nets learn to reconstruct an incomplete 3D shape by extracting feature pyramids that retain global and local shape geometry priors. To allow for dynamic reconstruction, some works complete a sequence of depth observations without computing correspondences over time, e.g. STIF [59]. 4DComplete [32] completes the geometry and estimates the motion from one partial geometry and motion field observation. Zhou et al. [60] complete the geometry and estimate the motion using two time frames containing partial geometric observations of a 3D shape. These works are limited to reconstructing sequences of one or two observations and do not benefit from long-range temporal information.
@@
In contrast, we do not decouple shape and motion. Instead, we reconstruct each frame using signed distances allowing for high fidelity reconstructions. The temporal consistency of these SDFs is constrained using a near-isometric deformation model. As a result, the reconstructions can have high levels of geometric detail while being precisely tracked. Table 1 positions our work w.r.t. competing methods according to their ability to provide a dense tracking, complete partial inputs, take into account long temporal context (more than 2 frames), preserve geometric detail present in the input, and train without inter-frame correspondence supervision. Our method is the only one that fulfills all five desiderata.
@@ 3 Method
Figure 2: Overview of our approach. Given Truncated Signed Distance Field grids representing partial observations of a moving 3D shape (leftmost), our approach achieves detailed reconstructions with dense tracking (rightmost). During Feature-Fusion Based Completion (purple module), the TSDFs are encoded in a latent space where self-attention allows to fuse and complete the observed information. The fused latent features are decoded into coarse shapes and then refined where this coarse surface locates. During Inter-Frame Deformation Estimation (green module), these reconstructions are fitted to a patch-wise near-rigid mesh deformation model that implements a near-isometric deformation assumption promoting their consistency.
Given partial observations of a moving 3D shape, we compute both complete shapes and their temporal evolution. To model the latter, recent methods either rely on a parametric shape model limiting generalisation to different shape classes, or on an unconstrained 3D flow that deforms a template shape leading to distortions and to a low preservation of observed geometric detail. Instead, we propose to use a near-isometric mesh deformation model. On the one hand, the near-isometric deformation assumption allows to represent a variety of moving shapes. On the other hand, mesh deformation modeling allows to control for the amount of surface distortions induced by the deformation, promoting the consistency of the reconstructions. Further, we do not decouple shape and motion. Rather, we propose a multi-scale feature-fusion strategy to represent each frame as a neural field able to capture observed geometric details, and then link these fields under the mesh deformation constraint. Fig. 2 gives an overview of our approach.
Our approach proceeds in two steps. First (Sec. 3.1) is a fusion and completion step, where the observations are encoded in a latent space, fused and then decoded into complete shapes. To efficiently produce high fidelity completions, we employ a multi-scale implicit surface representation using signed distances (purple module in Fig. 2). The fusion step trains with complete shape supervision to build a shape space for the completion task. Second, is the deformation search (Sec. 3.2) that translates the fusion and completion in feature space to a near-isometric deformation in 3D. This is achieved by fitting our implicit surfaces to a mesh-based near-isometric deformation model (green module in Fig. 2). Thanks to the near-isometric deformation assumption, our method does not require inter-frame correspondence to train, nor does it need a shape in a canonical pose for each sequence. Our method optimises for a fusion objective and for a self-supervised deformation objective:
@@
with fusion and deformation objectives $l^{fusion}$ and $l^{def}$, which are detailed in the following sections.
@@ 3.1 Feature-Fusion Based Completion
Given a sequence of $N$ TSDF (truncated signed distance field) volumes [14] $(\mathcal{T}_{i}\in\mathbb{R}^{D\times H\times W})_{i\in\{1,...,N\}}$ representing partial observations of a moving 3D shape, the feature-fusion based completion fuses these observations in a feature space to complete each frame while retaining observed geometric details. It computes a neural signed distance field $SDF_{\Theta}(i)$ for each frame $i$. To produce high fidelity reconstructions, we employ geometry aligned features [22, 32, 60] to represent $SDF_{\Theta}(i)$, that is, features that align with the underlying surface. To represent neural fields with high frequency geometric details, we propose a two-scale grid of geometry aligned features. We first extract a coarse grid denoted as $(\mathcal{F}^{c}_{i}\in\mathbb{R}^{D_{c}\times H_{c}\times W_{c}\times C})_{i\in\{1,...,N\}}$ representing coarse completions $SDF^{coarse}_{\Theta}(i)$. Then, to represent high frequency details, we only refine this grid where the coarse surface locates i.e. where $SDF^{coarse}_{\Theta}(i)$ is lower than a certain threshold. We denote the refined features as $(\mathcal{F}_{i}\in\mathbb{R}^{D_{F}\times H_{F}\times W_{F}\times C})_{i\in\{1,...,N\}}$.
To get the SDF value at query point $x$ for the completion of frame $i$, we first interpolate the feature volume $\mathcal{F}_{i}$ using trilinear interpolation, and pass the resulting feature concatenated with $x$ to an MLP denoted as $S_{\Sigma}$ to yield the desired value. Given that each $\mathcal{T}_{i}$ is normalised in a bounding box $B$, we can write $SDF_{\Theta}(i)$ as follows:
@@
with coarse-scale and fine-scale objectives $l^{coarse}$ and $l^{fine}$, which are detailed in the following sections.
The geometry aligned features are extracted using a feature extractor $F_{\Psi}$. It acts in four steps. First, $F_{\Psi}$ encodes each $\mathcal{T}_{i}$ in a latent feature space (Sec. 3.1.1). Then, using an attention mechanism, it fuses the latent features to share the observed information at each frame (Sec. 3.1.2). The fused latent features are decoded into coarse feature volumes representing a coarse completion of each frame (Sec. 3.1.3). In order to produce high fidelity reconstructions, the coarse feature volumes are refined only where the coarse surface locates, using features describing fine details retained during encoding (Sec. 3.1.4).
@@ 3.1.1 Frame-Wise Latent Feature Encoder
$F_{\Psi}$ first encodes the information observed in each $\mathcal{T}_{i}$ in a latent feature space. In the interest of efficiency, we leverage a sparse convolutional encoder [18] (i.e. convolutions are only applied at grid locations where $\mathcal{T}_{i}$ is defined). This brings us to a coarser spatial resolution $D_{c}\times H_{c}\times W_{c}$ where standard 3D convolution is computationally feasible. Then, a 3D convolutional encoder intervenes to yield our latent per-frame features $(\mathcal{F}^{latent}_{i}\in\mathbb{R}^{d_{l}})_{i\in\{1,...,N\}}$.
@@ 3.1.2 Feature Fusion
In order to allow for geometric fusion and completion, $F_{\Psi}$ communicates the information encoded in the latent features between frames. This is done thanks to a self-attention mechanism [52] applied on the latent codes $\mathcal{F}^{latent}_{i}$. This means that the latent features outputted by this self-attention mechanism that we denote $(\mathcal{F}^{latent^{1\rightarrow N}}_{i}\in\mathbb{R}^{d_{l}})_{i\in\{1,...,N\}}$ encode our fused and completed shape information.
@@ 3.1.3 Coarse-Dense Reconstruction
The latent features encoding fused and completed shapes need to be decoded into feature volumes able to capture observed geometric detail. Inspired by [32, 60], we propose a two-scale grid of features to define each neural field. We first locate the underlying surface at a coarse resolution, and then only refine features where this surface locates. The coarse-dense reconstruction step is responsible for locating this coarse surface. In the geometry aligned representation of neural fields, this translates to finding a coarse feature volume $(\mathcal{F}^{c}_{i}\in\mathbb{R}^{D_{c}\times H_{c}\times W_{c}\times C})_{i\in\{1,...,N\}}$. In the absence of information about where shape parts might be in our bounding box $B$, we must obtain dense features, i.e. features everywhere in $B$, so we can interpolate these features at any spatial location $x$ to obtain the associated SDF value. To that end, we employ a 3D convolutional decoder on our latent codes $(\mathcal{F}^{latent^{1\rightarrow N}}_{i})_{i\in\{1,...,N\}}$ to compute the coarse-dense feature volumes $(\mathcal{F}^{c}_{i})_{i\in\{1,...,N\}}$. As training objectives, we impose that the neural fields encoded by these features approximate our ground truth SDF values and that they retain the SDF property, using the following losses:
@@
$\footnotesize l^{SDF}_{coarse}=\frac{1}{N}\sum\limits_{i=1}^{N}\frac{1}{S}\sum\limits_{j=1}^{S}(|S_{\Sigma}(\operatorname{tri}(\mathcal{F}^{c}_{i},x_{j}),x_{j})-gt^{i}_{sdf}(x_{j})|)$ (5)
$\footnotesize l^{eikonal}_{coarse}=\frac{1}{N}\sum\limits_{i=1}^{N}\frac{1}{S}\sum\limits_{j=1}^{S}(\|\|\nabla_{x_{j}}S_{\Sigma}(\operatorname{tri}(\mathcal{F}^{c}_{i},x_{j}),x_{j})\|_{2}-1\|_{2}^{2})$ (6)
where $\lambda_{1},\lambda_{2}\in\mathbb{R}$ are weights for loss terms; $(x_{j}\in B)_{j\in\{1,...,S\}}$ are $S$ points sampled in $B$ and $gt^{i}_{sdf}(x_{j})$ is the ground truth SDF value at point $x_{j}$ for frame $i$

### v2

# NF3DM: Combining Neural Fields and Deformation Models for 3D Non-Rigid Motion Reconstruction

@@ 1 Introduction
Non-rigid 3D motion reconstruction involves recovering the shape and movement of objects undergoing arbitrary non-rigid motions based on possibly partial visual observations. Given our naturally dynamic world, this task has extensive applications, particularly in digitizing natural scenes for virtual reality and entertainment. Our focus is on reconstructing a moving 3D shape from a monocular depth video.
@@
Figure 1: Given monocular depth maps of a moving shape, our approach produces complete reconstructions that preserve observed geometric details while establishing dense tracking. We experiment with motions of clothed humans (left) and animals (right).
@@
The first one encompasses parametric models, which use combined shape and motion parameters to handle specific entities (e.g. humans, faces, or animals). Examples include models like SCAPE [4], BlendSCAPE [23], and SMPL [36] for human figures, Flame [32] for faces , and SMAL [67] for animals. These parametric models have gained significant success, largely thanks to their ability to provide robust, temporally consistent estimations. However, they lack generalizability across shape classes and struggle to capture geometric details outside of model constraints, such as hair or loose clothing in human representations.
@@
The second category includes methods that decouple shape and motion models, allowing for generalization across shape classes. Inspired by Niemeyer et al[44], approaches in this category use an implicit scene representation as a template, that undergoes unconstrained displacement fields (e.g., 3D flows) over time. While templates enforce consistency, they can restrict the model’s ability to depict finer geometric details. Additionally, 3D flow-based motion modelling strategies are insufficiently constrained, leading to distorted surfaces. Another branch in this category focuses on time-based shape reconstructions rather than motion, either omitting explicit motion modeling altogether [62] or limiting it to frame pairs [33, 63]. Although useful for shape capture, these approaches lack broader applicability where motion is required and may yield topologically inconsistent reconstructions.
@@
We propose a novel data-driven representation of 3D motions that combines an implicit neural field shape representation with a near-isometric mesh deformation model. This combination enables tracking a large class of 3D motions such as clothed 3D human motions and other vertebrate motions such as animals. As neural field representations allow for detailed reconstructions and flexible topologies, we leverage the near-isometric deformation to promote temporal coherence and minimal surface distortions, which helps resolve shape inconsistencies that may arise from independent reconstructions.
@@
To this end, we introduce a data-driven approach that combines two modules working in synergy. A fusion module, that fuses input observations over a time sequence to infer full neural field reconstructions at each time step using an encoder-decoder architecture equipped with an attention mechanism. And a deformation module, that predicts inter-frame deformations by fitting the neural field reconstructions to a near-isometric mesh deformation model using a deformation network. Both modules train together without motion supervision but with losses that promote geometric feature association between frames, near-isometric deformations, and 3D reconstruction losses. After training, per-frame reconstructions and their tracking are inferred from monocular observations in a single forward step.
@@
  • Investigate and demonstrate the benefits of combining neural field representations with explicit surface deformation models when handling temporal sequences.
@@ 2 Related Works
Methods to reconstruct a possibly moving 3D shape can be categorized into two classes. On the one hand, array-based methods [7, 59, 56, 65, 66, 34, 20, 24] use multiple calibrated cameras that require costly setups and are thus restricted to professional use. On the other hand, depth-fusion-like methods [25, 21, 51, 60, 15] enable data obtained from commodity sensors to be used for consumer level applications. An active research direction aims to reconstruct possibly moving 3D shapes from sensor data, e.g. from a single RGB or RGB-D image [50, 22, 48], from an RGB video [2, 17, 3, 31, 10], from depth views [9, 62, 63, 58] or from point clouds [44, 52]. We review methods that input geometry observations, i.e. depth views or point clouds, as our method considers similar inputs. These approaches follow two main lines of works: model-based methods that leverage parametric shape models, and model-free methods that generalize to multiple shape classes.
@@ 2.1 Model-Based Methods
Early model-based strategies propose optimisation-based techniques. Weiss et al[57] fit partial observations to the SCAPE [4] human body model. Mosh [35] uses a sparse set of markers to fit SCAPE [4] while Mosh++ [39] uses SMPL [36]. To augment the expressivity of parametric human models, several approaches [49, 1, 55] add vertex displacements on top of SMPL [36] to model clothes, while DoubleFusion [61] proposes to add a 3D graph. More recently, data-driven approaches were proposed. IP-Net [6] combines parametric and implicit representations. H4D [26] proposes a compositional representation, which disentangles shape and motion. NSF [58] propose to combine SMPL [36] with a neural surface field to represent fine grained surface details. Neural Parametric Models (NPMs) [45] propose to learn custom disentangled shape and pose spaces from a dataset to which observations can be fitted at inference. SPAMs [46] extend NPMs by learning disentangled semantic-part-based shape and pose spaces.
@@ 2.2 Model-Free Methods
Method
Tracking
Shape
Completion
Long
Temp. Ctxt.
Detail
Preservation
Unsupervised
@@
Some works consider static 3D reconstruction, e.g. Implicit Feature Networks (IF-Nets) [13]. IF-Nets learn to reconstruct an incomplete 3D shape by extracting feature pyramids that retain global and local shape geometry priors. To allow for dynamic reconstruction, some works complete a sequence of depth observations without computing correspondences over time, e.g. STIF [62]. 4DComplete [33] completes the geometry and estimates the motion from one partial geometry and motion field observation. Zhou et al[63] complete the geometry and estimate the motion using two time frames containing partial geometric observations of a 3D shape. These works are limited to reconstructing sequences of one or two observations and do not benefit from long-range temporal information.
@@
In contrast, our approach does not decouple shape and motion. Instead, each frame is reconstructed using signed distances, allowing for high fidelity reconstructions. The temporal consistency of these SDFs is constrained using a near-isometric deformation model. As a result, the reconstructions can have high levels of geometric detail while being precisely tracked. Table 1 positions our work w.r.t. competing methods according to their ability to provide a dense tracking, complete partial inputs, take into account long temporal context (more than 2 frames), preserve geometric detail present in the input, and train without inter-frame correspondence supervision. Our method is the only one that fulfills all five desiderata.
@@ 3 Method
Figure 2: Overview of our approach. Given Truncated Signed Distance Field grids representing partial observations of a moving 3D shape (leftmost), the method achieves detailed reconstructions with dense tracking (rightmost). During Feature-Fusion Based Completion (purple module), the TSDFs are encoded in a latent space where self-attention allows to fuse and complete the observed information. The fused latent features are decoded into coarse shapes and then refined where this coarse surface locates. During Inter-Frame Deformation Estimation (green module), the fusion is further constrained by fitting the reconstructions to a patch-wise near-rigid mesh deformation model that implements a near-isometric deformation assumption promoting their consistency.
Given partial observations of a moving 3D shape, our approach estimates both complete shapes and their temporal evolution. To model the latter, recent methods rely on a parametric shape model limiting generalisation to different shape classes, or on an unconstrained 3D flow that deforms a template leading to distortions and to a low preservation of observed geometric detail. Instead, we use a near-isometric mesh deformation model. On the one hand, the near-isometric deformation assumption allows to represent a variety of moving shapes. On the other hand, mesh deformation modeling allows to control the amount of surface distortions induced by the deformation, promoting consistency of the reconstructions. Furthermore, shape and motion are not decoupled. Instead, we propose a multi-scale feature-fusion strategy to represent each frame as a neural field capable of capturing observed geometric details, and link these neural fields under the mesh deformation constraint. Fig. 2 gives an overview of the method.
Our approach relies on the synergy of two modules. A feature-fusion and completion module (Sec. 3.1), where the observations are encoded in a latent space, fused and then decoded into complete shapes. To efficiently produce high fidelity completions, we use a multi-scale implicit surface representation using signed distances (purple module in Fig. 2). The fusion module trains with complete shape supervision to build a shape space for the completion task. The fusion is further constrained by learning a surface deformation prior with a deformation search module (Sec. 3.2) that translates the fusion and completion in feature space to a near-isometric deformation in 3D. Implicit surfaces are fitted to a mesh-based near-isometric deformation model (green module in Fig. 2). Both modules are inter-dependent benefiting one another. With the near-isometric deformation assumption, our method trains without inter-frame correspondences and without requiring a shape in a canonical pose per sequence. It optimises for a fusion objective and for a self-supervised deformation objective:
@@
with fusion and deformation objectives $l^{fusion}$ and $l^{def}$.
@@ 3.1 Feature-Fusion Based Completion
Given a sequence of $N$ TSDF (truncated signed distance field) volumes $(\mathcal{T}_{i}\in\mathbb{R}^{D\times H\times W})_{i\in\{1,...,N\}}$, which can be readily obtained from depth videos with known camera parameters [14, 42] and represent partial observations of a moving 3D shape, the feature-fusion based completion fuses these observations in a feature space to complete each frame while retaining observed geometric details. It computes a neural signed distance field $SDF_{\Theta}(i)$ for each frame $i$. To produce high fidelity reconstructions, $SDF_{\Theta}(i)$ is represented with geometry aligned features [22, 33, 63]. These features are obtained at two scale levels. First a coarse grid $(\mathcal{F}^{c}_{i}\in\mathbb{R}^{D_{c}\times H_{c}\times W_{c}\times C})_{i\in\{1,...,N\}}$ is extracted, that represents coarse completions $SDF^{coarse}_{\Theta}(i)$. Second, with the aim to capture high frequency details, this grid is refined only where the coarse surface locates, i.e. where $SDF^{coarse}_{\Theta}(i)$ is lower than a certain threshold. The refined features are denoted as $(\mathcal{F}_{i}\in\mathbb{R}^{D_{F}\times H_{F}\times W_{F}\times C})_{i\in\{1,...,N\}}$.
To obtain the SDF value at a query point $x$ for the completion of frame $i$, we first interpolate the feature volume $\mathcal{F}_{i}$ using trilinear interpolation. The resulting feature, concatenated with $x$, is then passed to a MLP, denoted as $S_{\Sigma}$, to produce the desired value. Given that each $\mathcal{T}_{i}$ is normalised in a bounding box $B$, we can express $SDF_{\Theta}(i)$ as follows:
@@
with coarse-scale and fine-scale objectives $l^{coarse}$ and $l^{fine}$.
The geometry aligned features are extracted using a feature extractor $F_{\Psi}$ acting in four steps. First, it encodes each $\mathcal{T}_{i}$ in a latent feature space (Sec. 3.1.1). Then, using an attention mechanism, it fuses the latent features to share the observed information at each frame (Sec. 3.1.2). The fused latent features are decoded into coarse feature volumes representing coarse completions (Sec. 3.1.3). To get high fidelity reconstructions, the coarse feature volumes are refined where the coarse surface locates, using features describing fine details retained during encoding (Sec. 3.1.4).
@@ 3.1.1 Frame-Wise Latent Feature Encoder
$F_{\Psi}$ first encodes the information observed in each $\mathcal{T}_{i}$ in a latent feature space. For efficiency, we leverage a sparse convolutional encoder [18] (i.e. convolutions are only applied at grid locations where $\mathcal{T}_{i}$ is defined). This reduces the feature grid to a coarser spatial resolution $D_{c}\times H_{c}\times W_{c}$ where standard 3D convolution is computationally feasible. A 3D convolutional encoder then processes this grid to produce the latent per-frame features $(\mathcal{F}^{latent}_{i}\in\mathbb{R}^{d_{l}})_{i\in\{1,...,N\}}$.
@@ 3.1.2 Feature Fusion
In order to allow for geometric fusion and completion, $F_{\Psi}$ communicates the information encoded in the latent features between frames. This is done thanks to a self-attention mechanism [54] applied on the latent codes $\mathcal{F}^{latent}_{i}$. This means that the latent features outputted by this self-attention mechanism that we denote $(\mathcal{F}^{latent^{1\rightarrow N}}_{i}\in\mathbb{R}^{d_{l}})_{i\in\{1,...,N\}}$ encode fused and completed shape information.
@@ 3.1.3 Coarse-Dense Reconstruction
The latent features encoding fused and completed shapes must be decoded into feature volumes capable of capturing observed geometric details. Inspired by [33, 63], we use a two-scale grid of features to define each neural field. The underlying surface is first located at a coarse resolution, and then features are refined where this surface locates. The coarse-dense reconstruction step locates this coarse surface. This translates to finding coarse feature volumes $(\mathcal{F}^{c}_{i}\in\mathbb{R}^{D_{c}\times H_{c}\times W_{c}\times C})_{i\in\{1,...,N\}}$. Since the locations of unobserved shape parts within the bounding box $B$, are unknown, we must obtain dense features, i.e. features present throughout $B$, allowing interpolation at any spatial location $x$ to retrieve the associated SDF value. To achieve this, we employ a 3D convolutional decoder on the latent codes $(\mathcal{F}^{latent^{1\rightarrow N}}_{i})_{i\in\{1,...,N\}}$ to compute the coarse-dense feature volumes $(\mathcal{F}^{c}_{i})_{i\in\{1,...,N\}}$. As training objectives, we enforce that the neural fields encoded by these features approximate the ground truth SDF values while preserving the SDF property, using the following losses:
@@
$l^{SDF}_{coarse}=\frac{1}{N}\sum\limits_{i=1}^{N}\frac{1}{S}\sum\limits_{j=1}^{S}(|S_{\Sigma}(\operatorname{tri}(\mathcal{F}^{c}_{i},x_{j}),x_{j})-gt^{i}_{sdf}(x_{j})|)$ (5)
$l^{eikonal}_{coarse}=\frac{1}{N}\sum\limits_{i=1}^{N}\frac{1}{S}\sum\limits_{j=1}^{S}(\|\|\nabla_{x_{j}}S_{\Sigma}(\operatorname{tri}(\mathcal{F}^{c}_{i},x_{j}),x_{j})\|_{2}-1\|_{2}^{2})$ (6)
where $\lambda_{1},\lambda_{2}\in\mathbb{R}$ are weights for loss terms; $(x_{j}\in B)_{j\in\{1,...,S\}}$ are $S$ points sampled in $B$ and $gt^{i}_{sdf}(x_{j})$ is the ground truth SDF value at point $x_{j}$ for frame $i$ .